] Carry out the following calculations and conversions by hand (showing all your working). (a) Compute the sum (1011010)2 + (11011)2 (without changing the base). (b) Compute the sum (135)8 + (2357)8 (without changing the base). (c) Convert (197)10 to octal. (d) Convert (20A5.76)12 to decimal.


1
Expert's answer
2021-10-12T14:55:52-0400

a)


"\def\arraystretch{1.5}
 \begin{array}{c} 
 & & 1 & 1 & & 1 \\
 1 & 0 & 1 & 1 & 0 & 1 & 0_2 \\
& & 1 & 1 & 0 & 1 & 1_2\\
 \hline
1 & 1 & 1 & 0 & 1 & 0 & 1_2\\
 
\end{array}"

b)


"\def\arraystretch{1.5}
 \begin{array}{c} 
& 1 & 1 & \\
 & 1 & 3 & 5_8 \\
 2 & 3 & 5 & 7_8\\
 \hline
 2 & 5 & 1 & 4_8\\
 
\end{array}"

c)


"197\div8=24 \ R5"

"24\div8=3\ R0"

"(197)_{10}=(305)_8"

d)


"(20A5.76)_{12}"

Integer part:

"2\cdot12^3+0\cdot12^2+10\cdot12^1+5\cdot12^0"

"=3456+0+120+5=3581"

Fraction part:


"7\cdot\dfrac{1}{12}+6\cdot\dfrac{1}{12^2}=\dfrac{15}{24}=0.625"

"(20A5.76)_{12}=(3581.625)_{10}"